# MAPF-GPT: Imitation Learning for Multi-Agent Pathfinding at Scale

### v3

@@ Introduction
The model was trained to replicate the behavior of the expert policy using cross-entropy loss (i.e., log-loss) via mini-batch stochastic gradient descent, optimized with AdamW (Loshchilov and Hutter 2019). The target label for this loss is the ground-truth action index provided by the expert policy, LaCAM. LaCAM is a centralized solver that builds a path for all agents during the whole episode, leveraging information about the full environment state. In contrast, the trainable model relies solely on the local observation $o$ of each agent $u$.
@@
After training, this policy allows actions to be sampled from it. While an alternative could be to pick the action with the highest probability, we use sampling, recognizing the decentralized nature of the policy. Sampling selects actions according to the probability distribution given by the policy:
@@
We used $2000$ warm-up iterations and cosine annealing (Loshchilov and Hutter 2017), with a gradient clipping value of $1.0$ and a weight decay parameter of $0.1$. The entire 1B dataset was utilized to train the 85M model, which underwent 1M iterations with a batch size of $512$, resulting in $15.625$ epochs based on the gradient accumulation steps, set at $16$. For training the 6M and 2M models, we used portions of the 1B dataset – 150M and 40M, respectively. Additional details about the parameters influencing the training process are provided in the Appendix.
@@
In the first series of experiments, we compare 3 variants of MAPF-GPT varying in the number of parameters in their neural networks (2M, 6M, 85M) with the state-of-the-art learnable MAPF solvers: DCC (Ma, Luo, and Pan 2021) and SCRIMP (Wang et al. 2023) [^1] Please note that a plethora of learnable MAPF solvers exists however only few of them are tailored to the studied setting: MAPF with non-disappearing agents.. We use the pre-trained weights for DCC and SCRIMP. These weights were obtained by the authors while training on the random maps. Additionally, we present the results of LaCAM (Okumura 2024), which served as the expert centralized solver for data collection. For evaluation we used Random, Mazes, Warehouse, MovingAI maps (the latter two are out-of-distribution for all learnable solvers). The details on the maps and problem instances are given in the Appendix.
The results are presented in Figure 3, where the success rate of all solvers is shown. Clearly, MAPF-GPTs significantly outperform DCC and SCRIMP across all scenarios, with the largest model, MAPF-GPT-85M, being the absolute leader. MAPF-GPT-6M shows slightly better results on the Random and Mazes maps, and comparable results on the Warehouse and MovingAI maps to the smallest MAPF-GPT-2M model.
Figure 4 shows the Sum-of-Costs (SoC) achieved by the solvers relative to SoC of LaCAM (the lower - the better). As can be seen, MAPF-GPTs again outperform the other approaches and their performance correlates with the number of model parameters. Interestingly, there are some rare cases on Random maps where DCC and MAPF-GPT-85M outperformed LaCAM in terms of SoC. The same situation is observed for MAPF-GPT-6M on Mazes maps.
Runtime
In this experiment, we compare the runtime efficiency of each approach. The results are presented in Figure 5; each data point indicates the average time spent deciding the next action for all agents. All MAPF-GPT models scale linearly with the increasing number of agents. The largest model, MAPF-GPT-85M, shows a slightly higher runtime than DCC and SCRIMP with up to 96 agents. However, beyond 128 agents, the runtime is better, as it depends linearly on the number of agents. Notably, the MAPF-GPT-2M and MAPF-GPT-6M models are more than 13 times faster than SCRIMP and 8 times faster than DCC for 192 agents setup.
Figure 5: Runtime comparison of MAPF-GPT, DCC, and SCRIMP models on the warehouse environment. The plot shows the average time required to decide the next action for all agents as the number of agents increases.
@@
In this experiment, we study how each part of the information influences the performance of the MAPF-GPT agent. To address this, we trained a 6M parameter model with certain pieces of information masked that were provided to the original model. We examine four different cases: if there is no goal information for all agents (noGoal), if there is no greedy action provided (noGA), if there is no action history (noAH), and if the agent is trained without cost-to-go information (still retaining information about obstacles). The results are presented in Table 1.
Dataset 6M noGoal noGA noAH noC2G
@@
As can be seen, the 6M model without masking shows better results on the Random and Puzzles maps. The model trained without goal information shows better results on the MovingAI maps, with $88.4\%$ compared to $82.0\%$ for the original model. This improvement is likely due to the large map sizes, where most conflicts arise during the agents’ movement toward their goals, making the exact goal coordinates less critical. Additionally, the functionality was compensated by the greedy action, which indicates the direction to the goal.
Surprisingly, the model trained without action history shows better performance on the Mazes and Warehouse instances. This suggests that action history is not crucial for behavioral cloning, as LaCAM does not rely on it. Despite these results, we argue for retaining movement history information in the agent’s observation, which could be essential for fine-tuning the model in the environment using RL technique.
Masking greedy action and cost-to-go information degrades the performance of the model on all the testing tasks, highlighting their crucial role in effective pathfinding and conflict resolution.
@@
In addition to evaluating MAPF-GPT on the MAPF instances it was trained on, we also assessed its performance in a Life-Long MAPF (LMAPF) setup. Unlike regular MAPF problems, in LMAPF, each agent receives a new goal location every time it reaches its current one. In this setup, the primary objective is throughput, which is defined as the average number of goals reached by all agents per time step.
We evaluated MAPF-GPT in both zero-shot and fine-tuned configurations. To fine-tune the model, we generated an additional dataset using Mazes maps. For expert data, we employed the RHCR approach (Li et al. 2021), as LaCAM is not well-suited for LMAPF. The dataset contains 90 million (observation, action) pairs. We used MAPF-GPT-6M for this experiment.
The results are presented in Table 2. Even the zero-shot model is able to compete with other existing learning-based approaches, such as Follower (Skrynnik et al. 2024b) and MATS-LP (Skrynnik et al. 2024c). Moreover, in all cases, fine-tuning improved the results of the MAPF-GPT-6M model.
@@
Random Mazes Warehouse MovingAI
MAPF-GPT-6M 1.497 0.908 1.113 2.840
MAPF-GPT-6M-t 1.507 1.087 1.270 2.994
RHCR 2.164 1.554 2.352 3.480
Follower 1.637 1.140 2.731 3.271
MATS-LP 1.674 1.125 1.701 3.320
@@ Conclusion
In this work, we studied the MAPF problem as a sequential decision-making task. We proposed an approach to derive an individual policy based on state-of-the-art machine learning techniques, specifically (supervised) imitation learning from expert data —- MAPF-GPT. To train MAPF-GPT, we created a comprehensive dataset of expert MAPF solutions, transformed these solutions into (observation, action) pairs, tokenized them, and trained several transformer models (with varying numbers of parameters) on this data. Empirically, we demonstrate that even with a quite moderate number of parameters, such as 2M, MAPF-GPT significantly outperforms modern learnable MAPF competitors across a wide range of setups. Our results provide a clear positive answer to the question, ”Is it possible to create a strong learnable MAPF solver purely through imitation learning?”. The limitations of our approach are discussed in the Appendix.
Appendix A Appendix
Appendix B Implementation Details
@@
Appendix C Hyperparameters and Training Details
@@
Appendix D Dataset: Technical Details
@@
The resultant 1B dataset contains 500 such files, requiring 258GB of disk space. While training the MAPF-GPT-6M and MAPF-GPT-2M models, we randomly selected 75 and 20 files from this dataset, resulting in 150M and 40M datasets, respectively. The full dataset was used for training MAPF-GPT-85M model. The sizes of the datasets were selected according to Chinchilla paper (Hoffmann et al. 2022), which empirically found a scaling law of 20 tokens per model parameter, demonstrating the efficiency of a 70-billion-parameter model trained on 1.4 trillion tokens compared to larger models trained on fewer tokens.
Appendix E Benchmark
For empirical evaluation we have utilized the evaluation tools provided by the POGEMA benchmark (Skrynnik et al. 2024a). Specifically, we used the following set of maps: Random, Mazes, Warehouse (Li et al. 2021), MovingAI (Stern et al. 2019) and Puzzles. Table 5 contains information about the number of agents, maps, and instances per map (seeds) used in each set.
Table 5: Details about different sets of maps from POGEMA benchmark.
@@ Appendix E Benchmark
Figure 8 showcases examples of maps from the POGEMA benchmark. The Mazes and Random map sets are generated using the built-in generators within POGEMA. The Puzzles maps are handcrafted. The Warehouse is a single map with limitations on possible start and goal positions, which has been used in previous LMAPF-related studies (Li et al. 2021; Skrynnik et al. 2024b). Start locations on this map can only be generated on the left or right sides of the map, where there are no obstacles, while goal locations can only be placed beyond or below the obstacles in the center of the map. These constraints prevent the generation of instances with more than 192 agents on this map.
The final evaluated set of maps is MovingAI, referred to as MovingAI-tiles in the POGEMA benchmark. While the original MovingAI dataset includes various types of maps, including random and maze maps, the MovingAI-tiles set in POGEMA contains only city maps. Additionally, due to the large size ($256\times 256$) of city maps, they were divided into 16 tiles, resulting in maps of $64\times 64$ size. An example of such scenario is presented in Figure 9.
Figure 9: Example of a city map from the MovingAI-tiles set in the POGEMA benchmark. The original city maps, with dimensions of $256\times 256$, were divided into 16 tiles, resulting in smaller $64\times 64$ maps.
@@ Appendix F Limitations
Appendix G Evaluation on Puzzles
@@
MAPF-GPT-85M $1.00$ p m 0.00 $11.97$ p m 1.32
MAPF-GPT-6M $1.00$ p m 0.00 $14.14$ p m 2.65
MAPF-GPT-2M $0.99$ p m 0.02 $18.28$ p m 5.42
DCC $0.91$ p m 0.04 $31.49$ p m 9.38
SCRIMP $0.94$ p m 0.03 $33.52$ p m 9.70
LaCAM $1.00$ p m 0.00 $10.96$ p m 1.15
@@
MAPF-GPT-85M $0.94$ p m 0.03 $42.08$ p m 12.22
MAPF-GPT-6M $0.94$ p m 0.03 $49.48$ p m 13.45
MAPF-GPT-2M $0.93$ p m 0.04 $60.22$ p m 15.45
DCC $0.75$ p m 0.07 $86.73$ p m 16.71
SCRIMP $0.82$ p m 0.06 $93.31$ p m 19.91
LaCAM $0.96$ p m 0.03 $34.89$ p m 11.37
@@
MAPF-GPT-85M $0.92$ p m 0.04 $99.93$ p m 22.59
MAPF-GPT-6M $0.88$ p m 0.05 $124.15$ p m 25.43
MAPF-GPT-2M $0.81$ p m 0.06 $136.50$ p m 27.19
DCC $0.51$ p m 0.08 $170.77$ p m 20.53
SCRIMP $0.70$ p m 0.07 $186.10$ p m 27.51
LaCAM $0.94$ p m 0.03 $63.02$ p m 17.97

### v5

@@ Introduction
The model was trained to replicate the behavior of the expert policy using cross-entropy loss (i.e., log-loss) via mini-batch stochastic gradient descent, optimized with AdamW (Loshchilov and Hutter 2019). The target label for this loss is a ground-truth action index provided by the expert policy, LaCAM. LaCAM is a centralized solver that builds a path for all agents during the whole episode, leveraging information about the full environment state. In contrast, the trainable model relies solely on a local observation $o$ of each agent $u$.
@@
Once trained, this policy enables the sampling of actions from it. While an alternative could be to pick the action with the highest probability, we use sampling, recognizing the decentralized nature of the policy. Sampling selects actions based on the probability distribution given by the policy:
@@
We used $2000$ warm-up iterations and cosine annealing (Loshchilov and Hutter 2017), with a gradient clipping value of $1.0$ and a weight decay parameter of $0.1$. The entire 1B dataset was used to train the 85M model, which underwent 1M iterations with a batch size of $512$, resulting in $15.625$ epochs based on the gradient accumulation steps, set at $16$. For training the 6M and 2M models, we used portions of the 1B dataset – 150M and 40M, respectively. Additional details about the parameters influencing the training process are provided in the Appendix B.
@@
In the first series of experiments, we compare 3 variants of MAPF-GPT varying in the number of parameters in their neural networks (2M, 6M, 85M) with the state-of-the-art learnable MAPF solvers: DCC (Ma, Luo, and Pan 2021) and SCRIMP (Wang et al. 2023) [^1] Please note that while a plethora of learnable MAPF solvers exist, only a few are tailored to the studied setting: MAPF with non-disappearing agents.. We use the pre-trained weights for DCC and SCRIMP. These weights were obtained by the authors while training on the random maps. Additionally, we present the results of LaCAM (Okumura 2024), which served as the expert centralized solver for data collection. For evaluation we used Random, Mazes, Warehouse, Cities-tiles maps (the latter two are out-of-distribution for all learnable solvers). All maps and instances utilized during the evaluation were taken from (Skrynnik et al. 2025). The details on the maps and problem instances are given in the Appendix D.
The results are presented in Figure 3, where the success rate of all solvers is shown [^2] Please note that the previous arXiv version of the paper, as well as the version accepted to AAAI’25, contains different results for SCRIMP. This is due to a technical error occurred when running SCRIMP on our test instances. Unfortunately, the error was identified by us after the AAAI’25 conference.. Clearly, all variants of MAPF-GPT outperform both DCC and SCRIMP on Random and Mazes maps. On Cities-tiles the success rate of MAPF-GPT-85M is better than the one of DCC and is on par with SCRIMP. On Warehouse the latter solver is superior to all others when the number of agents exceeds 128. A possible explanation is that SCRIMP utilizes the so-called value-based tie-breaking mechanism, allowing the agents to iteratively re-select the actions, and this mechanism turns out to pe particularly valuable to the Warehouse setup.
Figure 4 shows the Sum-of-Costs (SoC) achieved by the solvers relative to the SoC of LaCAM (the lower - the better). As can be seen, MAPF-GPTs outperform the other approaches and their performance correlates with the number of model parameters. Interestingly, there are some rare cases on Random maps where DCC and MAPF-GPT-85M outperformed LaCAM in terms of SoC. The same situation is observed for MAPF-GPT-6M on Mazes maps.
@@
In this experiment, we study how each part of the information influences the performance of the MAPF-GPT agent. To address this, we trained a 6M parameter model with certain pieces of information masked that were provided to the original model. We examine four different cases: if there is no goal information for all agents (noGoal), if there is no greedy action provided (noGA), if there is no action history (noAH), and if the agent is trained without cost-to-go information (still retaining information about obstacles). We used an additional type of map, Puzzles, for this experiment. These maps are quite small ($5\times 5$) and are specifically designed to assess the capability of algorithms in solving complex scenarios where agents need to execute cooperative actions. The results are presented in Table 1.
Scenario 6M noGoal noGA noAH noC2G
@@
As can be seen, the 6M model without masking shows better results on the Random and Puzzles maps. The model trained without goal information shows better results on the Cities-tiles maps, with $88.4\%$ compared to $82.0\%$ for the original model. This improvement is likely due to the large map sizes, where most conflicts arise during the agents’ movement toward their goals, making the exact goal coordinates less critical. Additionally, the functionality was compensated by the greedy action, which indicates the direction to the goal.
Surprisingly, the model trained without action history shows better performance on the Mazes and Warehouse instances. This suggests that action history is not crucial for behavioral cloning, as LaCAM does not rely on it. Despite these results, we argue for retaining movement history information in the agent’s observation, which could be essential for further fine-tuning (we leave this for future work).
Masking greedy action and cost-to-go information degrades the performance of the model on all testing tasks, highlighting their crucial role in effective pathfinding and conflict resolution.
@@
In addition to evaluating MAPF-GPT on the MAPF instances it was trained on, we also assessed its performance in the Life-Long MAPF (LMAPF) setup. Unlike regular MAPF problems, in LMAPF, each agent receives a new goal location every time it reaches its current one. In this setup, the primary objective is throughput, which is defined as the average number of goals reached by all agents per time step.
Scenario 6M 6M tuned RHCR Follower MATS-LP
Random 1.497 1.507 2.164 1.637 1.674
Mazes 0.908 1.087 1.554 1.140 1.125
Warehouse 1.113 1.270 2.352 2.731 1.701
Cities-tiles 2.840 2.994 3.480 3.271 3.320
@@
We evaluated MAPF-GPT in both zero-shot and fine-tuned configurations. To fine-tune the model, we generated an additional dataset using Mazes maps. For expert data, we employed the RHCR approach (Li et al. 2021), as LaCAM is not well-suited for LMAPF. The dataset contains 90 million observation-action pairs. We used MAPF-GPT-6M for this experiment.
The results are presented in Table 2. Even the zero-shot model is able to compete with other existing learning-based approaches, such as Follower (Skrynnik et al. 2024a) and MATS-LP (Skrynnik et al. 2024b). Moreover, in all cases, fine-tuning improved the results of MAPF-GPT-6M.
@@
Runtime
In this experiment, we compare the runtime of the considered solvers. The results are presented in Figure 5; each data point indicates the average time spent deciding the next action for all agents. All MAPF-GPT models scale linearly with the increasing number of agents. The largest model, MAPF-GPT-85M, shows a slightly higher runtime than DCC and SCRIMP on the instances with up to 96 agents. However, beyond 128 agents, the runtime of MAPG-GPT-85M is better, as it depends linearly on the number of agents. Notably, the MAPF-GPT-2M and MAPF-GPT-6M models are more than 13 times faster than SCRIMP and 8 times faster than DCC for 192 agents setup.
Figure 5: Runtime of MAPF-GPT, DCC, and SCRIMP models on the Warehouse map. The plot shows the average time required to decide the next action for all agents as the number of agents increases.
@@ Conclusion
In this work, we studied the MAPF problem as a sequential decision-making task. We proposed the approach to derive an individual policy based on state-of-the-art machine learning techniques, specifically (supervised) imitation learning from expert data – MAPF-GPT. To train MAPF-GPT, we created the comprehensive dataset of expert MAPF solutions, transformed these solutions into observation-action pairs, tokenized them, and trained several transformer models (with varying numbers of parameters) on this data. Empirically, we demonstrate that even with a quite moderate number of parameters, such as 2M, MAPF-GPT significantly outperforms modern learnable MAPF competitors across the wide range of setups. Our results provide a clear positive answer to the question, “Is it possible to create a strong learnable MAPF solver purely through imitation learning?”. The limitations of our approach are discussed in the Appendix E.
Appendix
Appendix A Implementation Details
@@
Appendix B Hyperparameters and Training Details
@@
Appendix C Dataset: Technical Details
@@
The resultant 1B dataset contains 500 such files, requiring 258GB of disk space. While training the MAPF-GPT-6M and MAPF-GPT-2M models, we randomly selected 75 and 20 files from this dataset, resulting in 150M and 40M datasets, respectively. The full dataset was used for training MAPF-GPT-85M model.
Appendix D Benchmark
For empirical evaluation we have utilized the evaluation tools provided by the POGEMA benchmark (Skrynnik et al. 2025). Specifically, we used the following set of maps: Random, Mazes, Warehouse (Li et al. 2021), Cities-tiles (Stern et al. 2019) and Puzzles. Table 5 contains information about the number of agents, maps, and instances per map (seeds) used in each set.
@@ Appendix D Benchmark
Table 5: Details about different sets of maps from POGEMA benchmark.
@@
Figure 8 showcases examples of maps from the POGEMA benchmark. The Mazes and Random map sets are generated using the built-in generators within POGEMA. The Puzzles maps are handcrafted. The Warehouse is a single map with limitations on possible start and goal positions, which has been used in previous LMAPF-related studies (Li et al. 2021; Skrynnik et al. 2024a). Start locations on this map can only be generated on the left or right sides of the map, where there are no obstacles, while goal locations can only be placed beyond or below the obstacles in the center of the map. These constraints prevent the generation of instances with more than 192 agents on this map.
The final evaluated set of maps is Cities-tiles. This set of maps, as well as the others, is taken from POGEMA benchmark, however, these maps are based on the ones presented in the MovingAI benchmark. While the original MovingAI dataset includes various types of maps, including random and maze maps, the Cities-tiles set in POGEMA contains only city maps. Additionally, due to the large size ($256\times 256$) of city maps, they were divided into 16 tiles, resulting in maps of $64\times 64$ size. An example of such scenario is presented in Figure 9.
Figure 9: Example of a map from the Cities-tiles set in the POGEMA benchmark. The original city maps, with dimensions of $256\times 256$, were divided into 16 tiles, resulting in smaller $64\times 64$ maps.
@@ Appendix E Limitations
Appendix F Evaluation on Puzzles
@@
MAPF-GPT-85M $1.00$ p m 0.00 $11.37$ p m 1.27
MAPF-GPT-6M $1.00$ p m 0.00 $12.69$ p m 2.28
MAPF-GPT-2M $0.99$ p m 0.02 $16.73$ p m 5.42
DCC $0.91$ p m 0.04 $31.45$ p m 9.43
SCRIMP $0.95$ p m 0.04 $28.07$ p m 9.08
LaCAM $1.00$ p m 0.00 $10.43$ p m 1.07
@@
MAPF-GPT-85M $0.99$ p m 0.01 $27.92$ p m 6.59
MAPF-GPT-6M $0.99$ p m 0.02 $36.23$ p m 9.73
MAPF-GPT-2M $0.96$ p m 0.03 $51.07$ p m 14.00
DCC $0.76$ p m 0.07 $79.92$ p m 14.35
SCRIMP $0.84$ p m 0.06 $73.97$ p m 18.81
LaCAM $1.00$ p m 0.00 $19.01$ p m 1.92
@@
MAPF-GPT-85M $0.99$ p m 0.02 $66.30$ p m 12.52
MAPF-GPT-6M $0.91$ p m 0.04 $101.35$ p m 21.01
MAPF-GPT-2M $0.89$ p m 0.05 $122.19$ p m 24.64
DCC $0.54$ p m 0.08 $169.11$ p m 22.41
SCRIMP $0.76$ p m 0.06 $131.54$ p m 23.82
LaCAM $1.00$ p m 0.00 $33.09$ p m 3.33